OpenStudy (anonymous):

what is: (2.4*10^-4)/(3.0*10^7) Please can you do this in dividing standard form with the working out.

OpenStudy (anonymous):

is that 2.4 to the 10 power

OpenStudy (anonymous):

?

OpenStudy (anonymous):

or 2.4 times 10?

OpenStudy (jdoe0001):

\(\bf \large{ \cfrac{2.4\times10^{-4}}{3.0\times 10^7}\\ \quad \\ \textit{keep in mind that }\quad a^{-n} = \cfrac{1}{a^n}\\ \quad \\ \cfrac{2.4\times10^{-4}}{3.0\times 10^7}\implies \cfrac{2.4\times \frac{1}{10^4}}{3.0\times 10^7}\implies \cfrac{\frac{2.4}{10^4}}{3.0\times 10^7}\implies\cfrac{\frac{2.4}{10^4}}{\frac{3.0\times 10^7}{1}}\\ \quad \\ \textit{recall that }\quad \cfrac{\frac{a}{b}}{\frac{c}{d}} \implies \cfrac{a}{b}\times\cfrac{d}{c}\\ \quad \\ \implies \cfrac{\frac{2.4}{10^4}}{\frac{3.0\times 10^7}{1}} \implies \cfrac{2.4}{10^4}\times \cfrac{1}{3.0\times 10^7} }\)

OpenStudy (anonymous):

^^^

OpenStudy (jdoe0001):

ryanvarghese1279 dunno, what does that give you?

OpenStudy (jdoe0001):

\(\bf \cfrac{2.4}{10^4}\times \cfrac{1}{3.0\times 10^7}\implies\cfrac{2.4}{3.0\cdot 10^7\cdot10^4} \)

OpenStudy (jdoe0001):

sooo... what would that give you?

OpenStudy (anonymous):

Using the way I have been taught at school I get 8*10 to the power of -4 However, the worksheet says it is 8 * 10 to the power of -12 which is right?

OpenStudy (jdoe0001):

recall the exponential rule of \(\Large a^n \times a^m = a^{n+m}\)

OpenStudy (anonymous):

I am doing standard form

OpenStudy (anonymous):

yes, doing all that I get my answer

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

so....

OpenStudy (jdoe0001):

\(\bf \cfrac{2.4}{10^4}\times \cfrac{1}{3.0\times 10^7}\implies\cfrac{2.4}{3.0\cdot 10^7\cdot10^4}\\ 2.4\cdot \cfrac{1}{3.0}\cdot \cfrac{1}{10^{11}}\implies 2.4\cdot 3^{-1}\cdot 10^{-11}\)

OpenStudy (jdoe0001):

well... 2.4 * 3 isn't 8... so... not sure on that one

OpenStudy (anonymous):

its divide 2.4 by 3.0. then change it t standard form which would be 8.0 * 10 to the power of -1

OpenStudy (jdoe0001):

ohhh I see.. yes, it's

OpenStudy (anonymous):

so... now what do i do? I need this urgently

OpenStudy (jdoe0001):

\(\bf 2.4\cdot \cfrac{1}{3.0}\cdot \cfrac{1}{10^{11}}\implies \cfrac{2.4}{3.0}\cdot \cfrac{1}{10^{11}}\implies 0.8\cdot 10^{-11}\\ \quad \\ 8\cdot10^{-1}\cdot10^{-11}\implies 8\cdot 10^{-12}\)

OpenStudy (anonymous):

what does "." mean decimal point?

OpenStudy (jdoe0001):

ohh... the dot notation? times

OpenStudy (anonymous):

k

OpenStudy (anonymous):

cause the question is divison; wouldn't you have to divide 2.4 by 3.0

OpenStudy (anonymous):

and 10^4-10^7?

OpenStudy (jdoe0001):

yes you would.... so... what part confuses you?

OpenStudy (anonymous):

the sudden move from dividing to multiplying

OpenStudy (jdoe0001):

you mean this part => \(\bf 2.4\cdot 3^{-1}\cdot 10^{-11}\)

OpenStudy (jdoe0001):

well, the lovely site, as usual lately, went down for a bit anyhow , so, yes, you're correct, that was a mistake of mine.... it'd have ended up in a division either way, they shouldn't multiply since the base is different... just wasn't looking but yes \(\bf 2.4\cdot 3^{-1}\cdot 10^{-11}\implies \cfrac{2.4}{3}\cdot 10^{-11}\)

OpenStudy (jdoe0001):

sorry it took me this long to reply... but I got extremely lagged, then the site went down for a few, and ended up doing an errand